OpenStudy (anonymous):

What is the sum? 8/17 + 5/17 A. 13/289 B. 40/289 C. 13/34 D. 13/17

OpenStudy (anonymous):

It is D

OpenStudy (anonymous):

Answer is 13/17

OpenStudy (anonymous):

Since they already have the same denominators, you can already add up the numerator. so \[\frac{ 8 }{ 17 }+\frac{ 5 }{ 17 }=\frac{ 13 }{ 17 }\]

OpenStudy (unklerhaukus):

\[\frac{a}{c}+\frac bc=\frac{a+b}c\]

OpenStudy (anonymous):

i already knew it was d